There were 678 tickets purchased for a major league baseball game. The lower box tickets costs $12.50 and the upper box tickets costs $10.00. The total amount of money spent was $7515.00. How many of each kind of ticket were purchased?
How many lower box tickets were purchased?
How many upper box tickets were purchased?

Answer by TimothyLamb(4379)

1250L + 1000U = 751500
L + U = 678
L = 678 - U
---
1250L + 1000U = 751500
1250(678 - U) + 1000U = 751500
1250(678 - U) + 1000U = 751500
847500 - 1250U + 1000U = 751500
96000 = 250U
Answer:
U = 384
L = 294
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php